OpenStudy (anonymous):

Determine g-1(x) of the function g(x) = 12x + 6 where x≠ -3. please help i have no idea how to do this

OpenStudy (anonymous):

\[y=12x+6\] solve for "x" i.e., get \[x=\text{something in "y"}\]

OpenStudy (anonymous):

i got 1/2-y=6x.... i dont think thats right

OpenStudy (anonymous):

\[y=12x+6\\ y-6=12x\\ {y-6\over12}=x\]

OpenStudy (anonymous):

get it now?

OpenStudy (anonymous):

ohhhhhh got it!

OpenStudy (anonymous):

but thats not an answer choice..

OpenStudy (anonymous):

yep. so, \[g^{-1}(x)={x-6\over12}\]

OpenStudy (anonymous):

can i show you my answer choices

OpenStudy (anonymous):

a)g-1(x) = 2x - 6, x≠0 B)g-1(x) = 12x - 3, x≠0 C)g-1(x) = 12x - 6, x≠0 D)g-1(x) = 12x - 16, x≠0

OpenStudy (anonymous):

ok.. seems to me like this is "RECIPROCAL" not INVERSE

OpenStudy (anonymous):

wait one sec

OpenStudy (anonymous):

it didnt show up right when i posted it

OpenStudy (anonymous):

it should look like this